### v2

# Data Augmentation for Supervised Graph Outlier Detection with Latent Diffusion Models

Abstract
Graph outlier detection is a prominent task of research and application in the realm of graph neural networks. It identifies the outlier nodes that exhibit deviation from the majority in the graph. One of the fundamental challenges confronting supervised graph outlier detection algorithms is the prevalent issue of class imbalance, where the scarcity of outlier instances compared to normal instances often results in suboptimal performance. Conventional methods mitigate the imbalance by reweighting instances in the estimation of the loss function, assigning higher weights to outliers and lower weights to inliers. Nonetheless, these strategies are prone to overfitting and underfitting, respectively. Recently, generative models, especially diffusion models, have demonstrated their efficacy in synthesizing high-fidelity images. Despite their extraordinary generation quality, their potential in data augmentation for supervised graph outlier detection remains largely underexplored.
To bridge this gap, we introduce GODM, a novel data augmentation for mitigating class imbalance in supervised Graph Outlier detection with latent Diffusion Models. Specifically, our proposed method consists of three key components: (1) Variantioanl Encoder maps the heterogeneous information inherent within the graph data into a unified latent space. (2) Graph Generator synthesizes graph data that are statistically similar to real outliers from latent space, and (3) Latent Diffusion Model learns the latent space distribution of real organic data by iterative denoising. Extensive experiments conducted on multiple datasets substantiate the effectiveness and efficiency of GODM. The case study further demonstrated the generation quality of our synthetic data. To foster accessibility and reproducibility, we encapsulate GODM into a plug-and-play package and release it at the Python Package Index (PyPI): https://pypi.org/project/godm/.
Index Terms:
Graph Outlier Detection, Data Augmentation, Diffusion Models, Class Imbalance, Anomaly Detection
I Introduction
Graphs are foundational data structures that capture complex relationships among diverse entities. Recently, graph data mining has showcased effectiveness and flexibility across various tasks, including community classification among academic citation networks [1], recommender systems in social networks [2], knowledge graph completion [3], and so on and forth. Within this broad spectrum of applications, graph outlier detection has emerged as a popular and important area of research and practice [4, 5, 6]. Graph outlier detection focuses on detecting outliers within graph-structured data that significantly deviate from standard patterns, which has proven valuable in domains such as financial fraud detection in bank systems [7], fake news detection on social media [8], spam detection on review platform [9], and anti-money laundering in transaction networks [10].
However, despite the advancements in graph outlier detection techniques, similar to supervised outlier detection on other data modalities, graph outlier detection suffers a fundamental challenge known as class imbalance, where the number of positive instances (outliers) is significantly lower than the number of negative instances (inliers). For example, the ratio of positive to negative is only 1:85 in the DGraph dataset [7], reflecting the real-world extreme ratio in financial fraud detection scenario. This class imbalance problem poses challenges in the training of outlier detectors and often results in suboptimal performance. Specifically, as the negative instances dominate the training data, the loss function can be biased towards the majority of the negative class and, hence, exhibit poor generalization capability in identifying true outliers. In the worst case, a detector could classify every instance as negative, displaying seemingly low loss value and high accuracy score, yet basically lacking any knowledge.
Figure 1: An toy example of data augmentation for class imbalance in graph outlier detection. Synthetic outliers help balance the class distribution in the training data, thereby improving the performance of graph outlier detection.
Common practices mitigate this imbalance by upsampling or downsampling [11]. Upsampling augments the minority positive class by simply replicating outliers, whereas downsampling reduces the size of the majority negative class by randomly dropping normal instances. However, while beneficial in certain contexts, these methods often introduce their own challenges, such as the risk of overfitting outliers in the case of upsampling or the loss of valuable training data through downsampling. Another common approach to alleviate class imbalance is through instance reweighting in the loss function estimation. It allocates greater weights to positive instances and assigns lesser weights to negative instances. This approach is mathematically equivalent to upsampling and downsampling, thereby having a similar tendency to overfit outliers or underfitting inliers. These challenges underscore the need for more sophisticated data augmentation methods that can generate some synthetic data to handle the class imbalance in graph outlier detection, as Fig. 1 shows.
In recent years, generative models, particularly diffusion models, have achieved significant advancements in synthesizing high-fidelity image data [12]. Diffusion models capture intricate data distributions and generate high-quality samples by gradually denoising the samples from a simple prior distribution (e.g., Gaussian distribution). On graph data, although some works have explored the possibility of diffusion on molecular graphs [13, 14], few studies have been conducted to apply diffusion models to tackle the class imbalance problem in the task of graph outlier detection. However, existing diffusion models can hardly be directly adapted to large-scale graph outlier detection applications due to the following non-trivial challenges: (1) Heterogeneity: Unlike relatively simple molecular graphs, the information contained in graphs for outlier detection can be highly heterogeneous, including high dimensional features and more than one type of edge, even temporal information. Existing diffusion models are primarily designed for monotypic information, exemplified by the RGB channels in image data. (2) Efficiency: In the application of graph outlier detection, e.g., financial fraud detection, the graphs are typically much larger than molecular graphs, up to millions, even billions scale. The prohibitive computational cost of diffusion models hinders their direct application in outlier detection on large graphs. (3) Condition: Different from the unconditional generation on molecular graphs, conditional generation is required to mitigate the problem of class imbalance. We only want to generate outliers rather than normal nodes in synthetic graphs.
To bridge this gap, we make the first attempt to leverage diffusion models in graph outlier detection and propose a data augmentation for mitigating class imbalance in supervised Graph Outlier detection with latent Diffusion Models (GODM). Our main idea is to generate outliers in graph space while conducting diffusion in latent space. To address heterogeneity, we propose Variantioanl Encoder to map the heterogeneous information inherent within the graph data into a unified latent space. In addition, Graph Generator synthesizes different types of information back to graph space from the latent embedding. To alleviate the efficiency problem, instead of direct diffusion in graph space, we only conduct diffusion in the latent space crafted by the variational encoder. Furthermore, we use negative sampling and graph clustering to reduce the computational cost. For diffusion models, we also adopt EDM [15] instead of commonly used DDPM [16] to facilitate the generation efficiency of Latent Diffusion Model. For the condition, we not only give a class label to the variational encoder to form node embedding with class information but also conduct conditional generation on both Latent Diffusion Model and Graph Generator. Finally, our heterogeneous, efficient, and conditional GODM can generate graphs with outliers that are integrated with the original real graph for the training of the downstream graph outlier detector. Importantly, GODM is model agnostic, providing researchers and practitioners with the flexibility to integrate it across various graph outlier detectors.
Our contributions in this paper are mainly as follows:
  • We make the first attempt to leverage diffusion models in graph outlier detection and propose to adopt diffusion models in the latent space and bridge the graph space and the latent space with a variational encoder and graph generator for heterogeneity.
  • We further design outlier-only generation conditioned on outlier labels and improve the efficiency of GODM by negative sampling, graph clustering, and EDM.
  • We conduct extensive experiments to demonstrate the effectiveness and efficiency of GODM. To foster accessibility, we make our code a plug-and-play package, which is publicly available at: https://pypi.org/project/godm/.
The remainder of this paper is organized as follows. Section II formally defines the notation and the problem we studied in this paper. Section III provides an overview of GODM. Section IV gives detailed experimental results and analyses. Section V briefly describes related works. We summarize the paper and discuss future work in Section VI.
II Preliminary
Figure 2: The architecture of proposed data augmentation method GODM. For a given graph $\mathcal{G}$, GODM first partitions the large graph by graph clustering algorithms and conducts negative sampling on edge, reducing the computational cost. Then, a GNN-based Variational Encoder maps each node in subgraphs into the latent space embedding $\mathbf{Z}^{0}$. A forward diffusion process $\mathbf{Z}^{0}\rightarrow\mathbf{Z}^{S}$ is applied in the latent space. Reverse denoising process $\hat{\mathbf{Z}}^{S}\rightarrow\hat{\mathbf{Z}}^{0}$ starts from the prior distribution $p(\hat{\mathbf{Z}}^{S})$ and generates samples $\hat{\mathbf{Z}}^{0}$ in latent space with denoising MLP conditioned on outlier labels $\hat{\mathbf{y}}$. Graph Generator also conditionally generate node features $\hat{\mathbf{X}}$, graph structures $\hat{\mathcal{E}}$, edge timestamps $\hat{\mathbf{t}}$, and edge types $\hat{\mathbf{p}}$ from latent embedding $\hat{\mathbf{Z}}^{0}$. By ensembling all necessary information, we obtain the synthetic graph $\hat{\mathcal{G}}$. Along with the real graph ${\mathcal{G}}$, $\hat{\mathcal{G}}$ will be used for the training of the downstream graph outlier detector $D$.
In this section, we establish the notation adopted in the subsequent sections and rigorously formulate the problem of data augmentation for addressing class imbalance in supervised graph outlier detection.
II-A Notation
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X},\mathbf{y},\mathbf{t},\mathbf{p})$ denotes a graph with $n$ nodes, where $\mathcal{V}=\{v_{i}\}_{i=1}^{n}$ is the set of nodes, and =$\mathcal{E}=\{e_{ij}\}$ represents the set of edges. Here, $e_{ij}=(v_{i},v_{j})$ denotes an edge between node $v_{i}$ and node $v_{j}$. The matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ contains $d$-dimensional feature vectors $\mathbf{x}_{i}$ for each node $v_{i}$.  $\mathbf{y}\in\{0,1\}^{n}$ represents the vector of the node label $y_{i}$ for each node $v_{i}$, where 00 denotes an inlier, while $1$ denotes an outlier. $\mathbf{t}=\{t_{ij}\}\in\mathbb{N}$ and $\mathbf{p}=\{p_{ij}\}\in\{1,\ldots,P\}$ are the optional non-negative integer edge timestamp vector and edge type vector, respectively, where $P$ is the number of edge types.
II-B Problem Formulation
In this paper, we focus on the task of graph outlier detection, which is formally defined as:
Definition 1 (Graph Outlier Detection)
Given a graph $\mathcal{G}$, graph outlier detection can be regarded as a binary classification task that learns a detector $D:v_{i}\rightarrow\{0,1\}$ that classifies every node in $\mathcal{G}$ to an inlier (0) or an outlier (1).
In the task of graph outlier detection, we aim to mitigate class imbalance by data augmentation:
Definition 2 (Data Augmentation for Class Imbalance)
In outlier detection, the number of inliers is far more than the number of outliers, i.e., |$|\{v_{i}\mid y_{i}=0\}|\gg|\{v_{i}\mid y_{i}=1\}|$. Therefore, we aim to learn a parameterized data augmentation model, which can generate realistic and diverse synthetic graph $\hat{\mathcal{G}}=(\hat{\mathcal{V}},\hat{\mathcal{E}},\hat{\mathbf{X}},\hat{\mathbf{y}},\hat{\mathbf{t}},\hat{\mathbf{p}})$, where $\hat{y_{i}}=1,\forall v_{i}\in\hat{\mathcal{V}}$. To alleviate the class imbalance in the training of the graph outlier detector $D$, we integrate the original graph $\mathcal{G}$ and the synthetic graph $\hat{\mathcal{G}}$ for the training of detector $D$.
III Methodology
In this section, we elaborate on the proposed data augmentation method GODM in detail. Fig. 2 shows the architecture of GODM. It starts by partitioning the input graph $\mathcal{G}$ with graph clustering algorithms for reducing memory consumption. Subsequently, each partitioned subgraph is encoded into a latent space representation denoted by $\mathbf{Z}^{0}$ through Variational Encoder based on graph neural networks (GNNs), followed by a forward diffusion $\mathbf{Z}^{0}\rightarrow\mathbf{Z}^{S}$ in the latent space, where $S$ is the maximum step in the diffusion model. The reverse denoising process $\hat{\mathbf{Z}}^{S}\rightarrow\hat{\mathbf{Z}}^{0}$ iteratively estimates the noise and generates the latent embedding $\hat{\mathbf{Z}}^{0}$ from a predefined prior distribution $p(\hat{\mathbf{Z}}^{S})$ using a denoising function, which is a multi-layer perceptron (MLP), conditioned on class labels $\hat{\mathbf{y}}$. Graph Generator employs the latent representations $\hat{\mathbf{Z}}^{0}$ to conditionally reconstruct the node features and graph structures alongside other information available in the original real graph. By ensembling generated information, we are able to obtain $\hat{\mathcal{G}}$ that contains the nodes that are statistically similar to the real outliers. This synthetic graph $\hat{\mathcal{G}}$, in concert with the original graph $\mathcal{G}$, is leveraged to enhance the training efficacy of the downstream graph outlier detection task.
In Section III-A and Section III-B, we introduce the design details of the Variational Encoder and Graph Generator, respectively. Then, we describe Latent Diffusion Model in Section III-C. We summarize the training and inference process in Section III-D and Section III-E, respectively. Finally, we analyze the algorithmic complexity in Section III-F.
III-A Variational Encoder
Graph data is inherently complex and heterogeneous, containing both node feature $\mathbf{X}$ and graph structure $\mathcal{E}$, sometimes also partial node class labels $\mathbf{y}$, temporal information $\mathbf{t}$, and different types of edges $\mathbf{p}$. However, current diffusion models are mainly designed for monotypic information (e.g., the magnitude of RGB channels in images). To bridge this gap, we adopt a GNN-based Variational Encoder $E:\mathcal{G}\rightarrow\mathbf{Z}$ to map different types of information to a unified latent space.
Node Feature. To encode the node feature into the latent space, we take the feature of each node as the initial embedding for the encoder, i.e., $\mathbf{h}_{i}^{0}=\mathbf{x}_{i}$.
Class Label. As we only generate outliers by conditional generation, it will be helpful to encode node labels $\mathbf{y}=\{y_{i}\}$. We add the class information to the initial node embedding:
$\mathbf{h}_{i}^{0}=\mathbf{x}_{i}+\mathbf{w}^{C}_{E}\cdot y_{i},$ (1)
where $\mathbf{w}^{C}_{E}\in\mathbb{R}^{d}$ is a linear transformation for class label.
Graph Structure. Graph neural networks have emerged as a profound architecture for learning graph data, efficaciously learning the node representations by encoding the graph topology and node feature simultaneously with message passing. In this paper, we take GraphSAGE [17] as an example. For each layer of GraphSAGE, each node updates its embedding by aggregating the message within the neighborhood:
$\mathbf{h}_{i}^{l}=\text{ACT}(\mathbf{W}^{l}\cdot\text{CAT}(\mathbf{h}_{i}^{l-1},\text{AGG}(\{\mathbf{m}_{ij}^{l},\forall e_{ij}\in\mathcal{E}\}),$ (2)
where $\mathbf{h}_{i}^{l}$ is the node embedding for node $v_{i}$ in the $l-$th layer, $\mathbf{m}_{ij}^{l}$ is the edge-wise message of edge $e_{ij}$, and ACT is a non-linear activation function. $\mathbf{W}^{l}\in\mathbb{R}^{d^{l}\times 2d^{l-1}}$ is the linear transformation matrix, where $d^{l}$ and $d^{l-1}$ are the hidden dimension of $l$-th layer and $(l-1)$-th layer, respectively. While CAT represents concatenation, AGG denotes the aggregation function (e.g., mean). In the vanilla GraphSAGE, an edge-wise message is typically the embedding of source node $v_{j}$ from the last layer:
$\mathbf{m}_{ij}^{l}=\mathbf{h}_{j}^{l-1}.$ (3)
Edge Type. However, sometimes, we have different types of edges in graphs, known as heterogeneous graphs or heterogeneous information networks [18, 19]. As different edge types can encapsulate various semantics, encoding edge types can be important for the downstream task. Therefore, we add the type information to the edge-wise message:
$\mathbf{m}_{ij}^{l}=\mathbf{h}_{j}^{l-1}+\mathbf{W}^{P}\cdot\phi(p_{ij}),$ (4)
where $\mathbf{W}^{P}\in\mathbb{R}^{d^{l-1}\times P}$ is a linear transformation for edge type, and $\phi$ is the one-hot encoding function.
Edge Time. In addition, temporal information is critical in time series applications [20]. When edge time is available, we are also able to encode the timestamp to the edge-wise message with trigonometric temporal embedding [21]:
$\mathbf{m}_{ij}^{l}=\mathbf{h}_{j}^{l-1}+\mathbf{W}^{T}\cdot TE_{(t_{ij},\cdot)},$ (5)
where $\mathbf{W}^{T}\in\mathbb{R}^{d^{l-1}\times d^{l-1}}$ is the linear transformation for edge time, and $TE$ is defined as:
$TE_{(t_{ij},2k)}=sin(t_{ij}/10000^{2k/d^{l-1}}),$ (6)
$TE_{(t_{ij},2k+1)}=cos(t_{ij}/10000^{2k/d^{l-1}}),$
where $k$ is each dimension of the temporal embedding.
By stacking multiple layers, GNNs are capable of encoding both node features and neighborhood graph structures into independent and identically distributed (i.i.d.) node embedding $\mathbf{Z}$, which can be further leveraged by Latent Diffusion Model. To ease the generation for Latent Diffusion Model, we use Variational Encoder, which outputs two matrixes:
$\bm{\mu}=\text{GNN}_{\bm{\mu}}(\text{GNN}_{\text{shared}}(\mathcal{G})),$ (7)
$\text{log}\bm{\sigma}=\text{GNN}_{\bm{\sigma}}(\text{GNN}_{\text{shared}}(\mathcal{G})),$
[…]
$\mathbf{Z}=\bm{\mu}+\bm{\sigma}\cdot\bm{\varepsilon},\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I}),$ (8)
where $\mathcal{N}(\bm{0},\bm{I})$ refers to a multivariate normal (or Gaussian) distribution with a mean of $\bm{0}$ and a covariance of $\bm{I}$, where $\bm{0}$ denotes a zero vector, and $\bm{I}$ represents a identity matrix.
III-B Graph Generator
Graph Generator $G:\hat{\mathbf{Z}}\rightarrow\hat{\mathcal{G}}$ take the opposite process of Variational Encoder, generating $\hat{\mathcal{G}}=(\hat{\mathcal{V}},\hat{\mathcal{E}},\hat{\mathbf{X}},\hat{\mathbf{y}},\hat{\mathbf{t}},\hat{\mathbf{p}})$ given the latent space embedding $\hat{\mathbf{Z}}$. Each row $\hat{\mathbf{z}}_{i}$ in the embedding $\hat{\mathbf{Z}}$ is corresponding to a generated node $\hat{v}_{i}$ in $\hat{\mathcal{V}}$.
Class Label. Recall that our objective is exclusively generating outliers with positive labels. Consequently, rather than generating class labels, we take the desired labels as an input for conditional generation.
$\hat{\mathbf{z}}_{i}^{C}=\hat{\mathbf{z}}_{i}+\mathbf{w}^{C}_{G}\cdot\hat{y}_{i},$ (9)
where $\hat{\mathbf{z}}_{i}$ is the $i$-th row of $\hat{\mathbf{Z}}$ and $\hat{\mathbf{z}}_{i}^{C}$ is embedding with class condition. $\mathbf{w}^{C}_{G}\in\mathbb{R}^{d^{L}}$ is the linear transformation for the class label, and $d^{L}$ is the output dimension of the $L$-th layer of the encoder, i.e., the latent embedding dimension of $\hat{\mathbf{z}}_{i}$.
Node Feature. To generate the node feature $\hat{\mathbf{X}}$ from the latent space embedding, for the $i$-th row of $\hat{\mathbf{X}}$, we take the embedding with class condition $\hat{\mathbf{z}}_{i}^{C}$ as input:
$\hat{\mathbf{x}}_{i}=\mathbf{W}^{F}_{G}\cdot\hat{\mathbf{z}}_{i}^{C},$ (10)
where $\mathbf{W}^{F}_{G}\in\mathbb{R}^{d\times d^{L}}$ is the linear transformation for node feature generaion.
Graph Structure. The generation of graph structures denoted by the set of inferred edges $\hat{\mathcal{E}}$ requires link prediction between all pairwise combinations of nodes within the graph. The predicted edge score $\hat{e}_{ij}$ can be formulated as:
$\hat{e}_{ij}=\text{sigmoid}(\mathbf{w}^{E}_{G}\cdot\text{CAT}(\hat{\mathbf{z}}_{i}^{C},\hat{\mathbf{z}}_{j}^{C})),\forall(\hat{v}_{i},\hat{v}_{j})\in\mathcal{\hat{V}}\times\mathcal{\hat{V}},$ (11)
where $\mathbf{w}^{E}_{G}\in\mathbb{R}^{2d^{L}}$ is the linearn transformation for the edge generation. The generated edges are determined as follows:
^=$\hat{\mathcal{E}}=\{(\hat{v}_{i},\hat{v}_{j})\mid\hat{e}_{ij}\geq 0.5\}$ (12)
Edge Type. With the generated edges in hand, we can predict the type of every generated edge:
[…]
where $\mathbf{W}^{P}_{G}\in\mathbb{R}^{P\times 2d^{L}}$ is the linearn transformation for the edge type prediction.
Edge Time. Similarly, the timestamp of generated edges can be predicted as:
[…]
where $\mathbf{w}^{T}_{G}\in\mathbb{R}^{2d^{L}}$ is the linear transformation for the edge timestamp regression.
III-C Latent Diffusion Model
Utilizing Variational Encoder and Graph Generator, synthetic graph generation is already feasible. Nonetheless, the inherent complexity and heterogeneity of graph data pose a significant challenge when attempting a one-step estimation from a simple prior distribution, such as a Gaussian distribution, to the intricate target distribution. This challenge often leads to compromised generation quality and results in a marginal effect on the downstream graph outlier detection. Therefore, we integrate Latent Diffusion Model [12] to break down the estimation into a sequence of incremental steps. In each step, Latent Diffusion Model incrementally refines the distribution estimated, bridging the divergence between the simple prior and the intricate target distribution.
Latent Diffusion Model consists of a pair of processes. A fixed forward diffusion process perturbs the original data by the incremental adding Gaussian noise across a range of variances, and a reverse denoising process employs a learned denoising function to iteratively denoise the samples from a simple prior distribution, which is pure noise. In order to speed up the generation process, we adopt EDM [15] as our diffusion model. For the detailed design, we follow [22].
Forward Diffusion Process. We construct a forward diffusion process $\{\mathbf{Z}(s)\}_{s=0}^{S}$, where $s\in[0,S]$ is a continuous diffusion step. In the diffusion process, $\mathbf{Z}(0)=\mathbf{Z}$ is the embedding from Variational Encoder, while $\mathbf{Z}(S)\sim\mathcal{N}(\bm{0},\bm{I})$ is sampled from the prior distribution. According to [23], the forward diffusion process can be written in stochastic differential equation (SDE) as:
${\rm d}\mathbf{Z}=\bm{f}(\mathbf{Z},s){\rm d}s+g(s){\rm d}\bm{\omega}_{t},$ (15)
where $\bm{\omega}_{t}$ is the standard Wiener process (i.e., Brownian motion). $\bm{f}(\cdot,s)$ and $g(\cdot)$ are the drift coefficient and the diffusion coefficient, respectively. The selection of $\bm{f}(\cdot,s)$ and $g(\cdot)$ vary between different diffusion models. $f(\cdot)$ is usually of the form $\bm{f}(\mathbf{Z},s)=f(s)\mathbf{Z}$. Thus, Eq. 15 can be written as:
${\rm d}\mathbf{Z}=f(s)\mathbf{Z}{\rm d}s+g(s){\rm d}\bm{\omega}_{t}.$ (16)
Let $\mathbf{Z}$ be a function of diffusion step $s$, i.e., $\mathbf{Z}^{s}=\mathbf{Z}(s)$. The diffusion kernel of Eq. 16 can be represented in the conditional distribution of $\mathbf{Z}^{s}$ given $\mathbf{Z}^{0}$:
$p(\mathbf{Z}^{s}|\mathbf{Z}^{0})=\mathcal{N}(a(s)\mathbf{Z}^{0},a^{2}(s)\sigma^{2}(s)\bm{I}),$ (17)
where $a(s)$ and $\sigma(s)$ can be derived as:
$a(s)=\exp{\int_{0}^{s}f(\xi)\rm d\xi},$ (18)
$\sigma(s)=\sqrt{\int_{0}^{s}g^{2}(\xi)/a^{2}(\xi)\rm d\xi}.$ (19)
Consequently, the formulation of the forward diffusion process is equivalent to the definition of diffusion kernels characterized by $a(s)$ and $\sigma(s)$.
Conventional denoising diffusion probabilistic models (DDPM) [16] can be seen as discretizations of the variance preserving SDE with $a(s)=\sqrt{1-\beta(s)}$ and $\sigma(s)=\sqrt{\beta(s)/{(1-\beta(s))}}$, as $a^{2}(s)+a^{2}(s)\sigma^{2}(s)=1$. However, in order to achieve more efficient generation in GODM, we adopt EDM [15], which belongs to variance exploding SDE. Variance exploding SDE set $a(s)=1$, which implies that noise is directly added to the data instead of being blended through weighting. In this case, the variance of the noise (the noise level) is exclusively determined by $\sigma(s)$. In EDM, a linear noise level $\sigma(s)=s$ is applied. Therefore, the diffusion kernel can be written as:
$p(\mathbf{Z}^{s}|\mathbf{Z}^{0})=\mathcal{N}(\bm{0},\sigma^{2}(s)\bm{I}),$ (20)
and the forward diffusion process can be formulated as:
$\mathbf{Z}^{s}=\mathbf{Z}^{0}+\sigma(s)\bm{\varepsilon},\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I}).$ (21)
Reverse Denoising Process. As derived in [23], the reverse denoising process is formulated as the reverse SDE of Eq. 15:
${\rm d}\mathbf{Z}=[\bm{f}(\mathbf{Z},s)-g^{2}(s)\nabla_{\mathbf{Z}}\log p_{s}(\mathbf{Z})]{\rm d}s+g(s){\rm d}\bm{\omega}_{s},$ (22)
where $\nabla_{\mathbf{Z}}\log p_{s}(\mathbf{Z})$ is the score function of $\mathbf{Z}$. In EDM, given $a(s)=1$ and Eq. 18, it follows that,
$\bm{f}(\mathbf{Z},s)=f(s)\mathbf{Z}=\bm{0}.$ (23)
For the $g(s)$, derived from Eq. 19, we obtain
$g(s)=\sqrt{2\sigma(s)\dot{\sigma}(s)},$ (24)
where $\dot{\sigma}$ denotes the first order derivative of $\sigma$. With $\bm{f}(\mathbf{Z},s)$ and $g(s)$, we are able to obtain:
${\rm d}\mathbf{Z}=-2\dot{\sigma}(s)\sigma(s)\nabla_{\mathbf{Z}}\log p_{s}(\mathbf{Z}){\rm d}s+\sqrt{2\dot{\sigma}(s)\sigma(s)}{\rm d}\bm{\omega}_{s},$ (25)
where the noise level $\sigma(s)=s$.
III-D Training
Conventional graph generative models are highly constrained in scalability. Typical graph generative models are working molecular graphs that are at hundreds node scale [24]. However, outlier detection graphs are usually much larger, scaling to millions, even billions. Efficient training of generative models on large graphs requires special designs. We propose to apply negative sampling and graph clustering to improve the scalability of GODM.
Negative Sampling. In the graph structure generation, if the training of the generator includes every pair of nodes, this leads to a computational complexity of $O(n^{2})$. This parabolic complexity is catastrophic for large graphs (e.g., millions scale). Furthermore, it will result in a high imbalance in the training of the edge predictor itself. Negative sampling emerges as a crucial technique to reduce computational cost and alleviate the imbalance. To form a concise training set $\bar{\mathcal{E}}$ for edge generator, apart from adding the positive edge, i.e., $\{\bar{e}_{ij}=1\mid e_{ij}\in\mathcal{E}\}$, we randomly select a negative edge (a pair of nodes that is not connected) corresponding to every positive edge, $\{\bar{e}_{\tilde{i}\tilde{j}}=0\mid e_{\tilde{i}\tilde{j}}\notin\mathcal{E}\}$. By this means, we reduce the complexity from $O(n^{2})$ to $O(|\mathcal{E}|)$.
Graph Clustering. When the graph scales to millions of nodes, full-batch training becomes impractical, even with negative sampling. In addition, traditional neighbor sampling methods are well-suited in our case, as we need to reconstruct both the node feature and graph structure (i.e., edges). Specifically, node sampling cannot gather complete information for edge prediction, while edge sampling favors the nodes with a high degree. To address these challenges, we resort to graph clustering inspired by [25]. We first apply graph clustering algorithms (e.g., Metis [26]) to divide the large graph into small partitions. Then, we treat each partition as a minibatch and train GODM on each partition instead of the full graph.
Leveraging these two techniques, we are able to efficiently train GODM in two steps. We first train Variational Encoder and Graph Generator to bridge the graph space and latent space. Then, we train the diffusion model in the latent space.
Variational Encoder and Graph Generator. By integrating Variational Encoder and Graph Generator, we are able to train both of them in a variational autoencoder (VAE) fashion. We reconstruct the node feature with mean squared error:
$\ell_{\mathbf{X}}=\frac{1}{|\mathcal{V}|}\sum_{v_{i}\in\mathcal{V}}\|\mathbf{x}_{i}-\hat{\mathbf{x}}_{i}\|_{2}^{2},$ (26)
where $\|\cdot\|_{2}$ denotes the L2 norm. Then, we predict the edges in the training edge set $\bar{\mathcal{E}}$ with binary cross entropy loss:
$\ell_{\mathcal{E}}=-\frac{1}{|\bar{\mathcal{E}}|}\sum_{\bar{e}_{ij}\in\bar{\mathcal{E}}}(\bar{e}_{ij}\log(\hat{e}_{ij})+(1-\bar{e}_{ij}\log(1-\hat{e}_{ij}))).$ (27)
Similarly, we reconstruct the timestamp with mean squared error and edge type with cross-entropy loss:
$\ell_{\mathbf{t}}=\frac{1}{|\mathcal{E}|}\sum_{e_{ij}\in\mathcal{E}}(t_{ij}-\hat{t}_{ij})^{2},$ (28)
$\ell_{\mathbf{p}}=-\frac{1}{|\mathcal{E}|}\sum_{e_{ij}\in\mathcal{E}}\log(\hat{p}_{ij,p_{ij}}),$ (29)
We can obtain the total reconstruction loss by the weighted sum of all reconstruction loss:
$\ell_{\rm recon}=\omega_{\mathbf{X}}\ell_{\mathbf{X}}+\omega_{\mathcal{E}}\ell_{\mathcal{E}}+\omega_{\mathbf{t}}\ell_{\mathbf{t}}+\omega_{\mathbf{p}}\ell_{\mathbf{p}},$ (30)
where $\omega_{\mathbf{X}}$, $\omega_{\mathcal{E}}$, $\omega_{\mathbf{t}}$, and $\omega_{\mathbf{p}}$ are the hyperparameter weights balancing the importance between each term of reconstruction loss. Finally, we use the ELBO loss to train Variational Encoder and Graph Generator:
$\mathcal{L}_{\text{VAE}}=\ell_{\rm recon}+\beta\ell_{\rm kl}.$ (31)
where the $\ell_{\rm kl}$ is the Kullback-Leibler (KL) divergence loss between the latent space embedding and prior distribution (e.g., Gaussian distribution). However, as we have an additional Latent Diffusion Model, we loosen this regularization on the latent embedding. Consequently, we use a small hyperparameter weight $\beta<1$ to encourage the model to minimize reconstruction error while ensuring that the resultant embedding shape remains within the desired shape. The training of the VAE, consisting of Variational Encoder and Graph Generator, is presented in Algorithm 1.
Algorithm 1 GODM: Training of VAE
0:  Input graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X},\mathbf{y},\mathbf{t},\mathbf{p})$
0:  Node representation $\mathbf{Z}$, well-trained Variational Encoder $E$ and Graph Generator $G$
1:  Partition graph $\mathcal{G}$ into a set of subgraphs $\{\mathcal{G}_{\text{sub}}\}$
2:  for each $\mathcal{G}_{\text{sub}}$ do
[…]
4:     Reparameterization of $\mathbf{Z}$ via Eq. 8
5:     Generate $\hat{\mathbf{X}}$, $\hat{\mathcal{E}}$, $\hat{\mathbf{t}}$, $\hat{\mathbf{p}}$ via Eq. 10, 12, 14, 13
6:     Calculate loss via Eq. 31
7:     Update Variational Encoder and Graph Generator parameters via the Adam optimizer
8:  end for
Latent Diffusion Model. After the training of Variational Encoder and Graph Generator, we train Latent Diffusion Model via denoising score matching.
To solve the SDE in Eq. 25, we need to obtain the score function $\nabla_{\mathbf{Z}}\log p_{s}(\mathbf{Z})$, which is as intractable as $p_{s}(\mathbf{Z})=p(\mathbf{Z}^{s})$. However, the conditional distribution $p(\mathbf{Z}^{s}|\mathbf{Z}^{0})$ is tractable. From Eq. 17, we obtain its analytical solution:
$\nabla_{\mathbf{Z}}\log p(\mathbf{Z}^{s}|\mathbf{Z}^{0})=-\frac{\bm{\varepsilon}}{a(s)\sigma(s)}.$ (32)
Thus, we train a multi-layer perceptron (MLP) as the denoising function denoted by $\bm{\epsilon}_{\theta}$ for denoising score matching:
$\min\mathbb{E}_{\mathbf{Z}^{0}\sim p(\mathbf{Z}^{0})}\mathbb{E}_{\mathbf{Z}^{s}\sim p(\mathbf{Z}^{s}|\mathbf{Z}^{0})}\|\bm{\epsilon}_{\theta}({\mathbf{Z}}^{s}_{C},s)-{\bm{\varepsilon}}\|_{2}^{2},$ (33)
where ${\mathbf{Z}}^{s}_{C}=\mathbf{Z}^{s}+\mathbf{y}\mathbf{W}^{C},\mathbf{W}^{C}\in\mathbb{R}^{1\times d^{L}}$, encoding class label for conditional generation. The entire training process of Latent Diffusion Model is summarized in Algorithm 2.
Algorithm 2 GODM: Training of Latent Diffusion Model
0:  Node representation $\mathbf{Z}$
0:  Well-trained denoising function $\bm{\epsilon}_{\theta}$
1:  Sample the embedding $\mathbf{Z}^{0}$ from $p(\mathbf{Z})=p(\bm{\mu})$
2:  Sample diffusion steps $s$ from $p(s)$ then get $\sigma(s)$
3:  Sample noise vectors ${\bm{\varepsilon}}\sim\mathcal{N}(\bm{0},\sigma^{2}(s)\bm{I})$
4:  Get perturbed data ${\mathbf{Z}}^{s}=\mathbf{Z}^{0}+{\bm{\varepsilon}}$
5:  Calculate loss $\ell(\theta)=\|\bm{\epsilon}_{\theta}({\mathbf{Z}}^{s}_{C},s)-\bm{\varepsilon}\|_{2}^{2}$
6:  Update the network parameter $\theta$ via Adam optimizer
III-E Inference
The well-trained GODM is utilized in the inference process, augmenting real organic data by generated data. We first sampling $\hat{\mathbf{Z}}^{{S}}$ from Gaussian distribution. Then, we iteratively denoise $\hat{\mathbf{Z}}^{{s_{i}}}$ to obtain $\hat{\mathbf{Z}}^{{s_{i-1}}}$ conditioned on $\hat{\mathbf{y}}=\bm{1}$ via denoising function $\bm{\epsilon}_{\theta}$. The estimated $\hat{\mathbf{Z}}^{{0}}$ is fed into Graph Generator along with $\hat{\mathbf{y}}$ to generate synthetic graph $\hat{\mathcal{G}}$. Finally, the augmented graph $\mathcal{G}_{\text{sub}}$, which integrates the synthetic graph $\hat{\mathcal{G}}$ and the real organic graph $\mathcal{G}$, is used for training the downstream outlier detector $D$. Algorithm 3 presents the whole procedure of inference.
Algorithm 3 GODM: Inference
0:  Graph generator $G$ and denoising function $\bm{\epsilon}_{\theta}$
0:  Augmented graph $\mathcal{G}_{\text{aug}}$
1:  Sample $\hat{\mathbf{Z}}^{{S}}\sim\mathcal{N}(\bm{0},\sigma^{2}({S})\bm{I})$
2:  for $i=\max,\cdots,1$ do
3:     $\nabla_{\hat{\mathbf{Z}}^{s_{i}}}\log p(\hat{\mathbf{Z}}^{s_{i}})=-\bm{\epsilon}_{\theta}(\hat{\mathbf{Z}}^{s_{i}}_{C},{s_{i}})/\sigma({s_{i}})$
4:     get $\hat{\mathbf{Z}}^{s_{i-1}}$ via solving the SDE in Eq. 25
5:  end for
6:  Generate $\hat{\mathcal{G}}=G(\hat{\mathbf{Z}}^{0},\hat{\mathbf{y}})$
7:  $\mathcal{G}_{\text{aug}}=\text{batch}(\mathcal{G},\hat{\mathcal{G}})$
III-F Complexity Analysis
In this section, we meticulously analyze and elucidate the time efficiency and memory scalability of GODM in training and inference for both VAE (Variational Encoder and Graph Generator) and Latent Diffusion Model, providing insights for handling large graphs which are common in real-world applications of graph outlier detection.
III-F1 Time Efficiency
In the training of VAE, the major bottleneck comes from the graph structure generation in Graph Generator. This generation originally requires link prediction for every pair of nodes in the graphs, which is $O(n^{2})$. By graph clustering, we divide the large graph into relatively small partitions for minibatch training. We denote the average partition size (i.e., batch size) by $b$, so the number of partitions is $\frac{n}{b}$. Then, the complexity is reduced to $O(\frac{n}{b}b^{2})=O(nb)$, as only node pairs within each partition are predicted. We apply the negative sampling to further reduce the complexity to $O(|\mathcal{E}|)$, as we only sample a negative edge for each positive edge. For Latent Diffusion Model, we only sample a fixed number of diffusion steps to train the denoising function. As a result, assuming the number of latent dimensions is a constant, the complexity is $O(n)$. Thus, the total training complexity is $O(n+|\mathcal{E}|)$. For the inference, we first take $S$ steps sampling in Latent Diffusion Model, which is $O(nS)$, and predict edges between node pairs in each partition, which is $O(nb)$. Therefore, the total inference complexity is $O(nS+nb)$.
III-F2 Memory Scalability
The vanilla VAE requires memory consumption of $O(n^{2})$, which is infeasible for large graphs. To reduce memory consumption, we adopt minibatch training on graph partitions and negative sampling for link prediction. The complexity of VAE training can be obtained by the number of edges divided by the number of partitions $O(|\mathcal{E}|/\frac{n}{b})=O(\frac{|\mathcal{E}|b}{n})$. Latent Diffusion Model only requires a memory of $O(b)$. In the inference process, the VAE performs link prediction for all node pairs in a batch, leading to a memory complexity of $O(b^{2})$, and Latent Diffusion Model only demands $O(b)$ for each batch. Therefore, the total inference process needs $O(b^{2})$ of memory.
IV Experiments
In this section, we systematically delineate the experiments to evaluate GODM. We first introduce experimental setups in Section IV-A. Then, we aim to answer: RQ1 (Section IV-B): How effective is GODM on improving graph outlier detection performance? RQ2 (Section IV-C): How is the quality of the synthetic data generated by GODM? RQ3 (Section IV-D): How efficient is GODM in terms of time and memory? The code implementation of our experiments is publicly available at https://github.com/kayzliu/godm.
IV-A Experimental Setups
To benchmark GODM and contemporary graph outlier detection algorithms from various aspects, we conduct experiments in a unified and comprehensive environment, which is in part adapted from GADBench [6].
IV-A1 Datasets
TABLE I: Statistics of datasets
(*: with multiple edge types and temporal infomation).
Dataset #Nodes #Edges #Feat. Outlier
Weibo 8,405 407,963 400 10.3%
Tolokers 11,758 519,000 10 21.8%
Questions 48,921 153,540 301 3.0%
Elliptic 203,769 234,355 166 9.8%
DGraph 3,700,550 4,300,999 17 1.3%
We use five diverse datasets from GADBench. Weibo, Elliptic, Tolokers, and Questions are homogeneous static graphs with only one type of edge. DGraph is a heterogeneous temporal graph with different types of edges. The graphs vary in scale, ranging from thousands of nodes to millions of nodes. Table I provides the statistics of the datasets. In the table, #Nodes stands for the number of nodes, and #Edges stands for the number of edges. #Feat. denotes the raw feature dimension, i.e., the number of node attributes. The Outlier column represents the outlier ratio in the label, indicating the extent of class imbalance. The detailed descriptions for each dataset are as follows:
Weibo [27]: This dataset involves a graph delineating the associations between users and their corresponding hashtags from the Tecent-Weibo platform, consisting of 8,405 users and a collection of 61,964 hashtags. Activities within this dataset are regarded as suspicious (i.e., outliers) if they consist of a pair of posts occurring within narrowly defined temporal intervals (e.g., 60 seconds). Users engaging in at least five incidents of such behavior are classified under the suspicious category, in contrast to the remainder designated as benign. Following this classification criterion, the dataset contains 868 users identified as suspicious and 7,537 as benign. The primary feature vector includes geolocation data for each micro-blog entry and a representation utilizing the bag-of-words model.
Tolokers [28]: The dataset is obtained from the Toloka crowdsourcing platform. It is composed of nodes corresponding to individual workers who have engaged in at least one out of thirteen specified projects. Edges are established between pairs of workers who have concurrently contributed to an identical task. The primary objective is to predict the likelihood of a worker having received a ban in any one of the projects. The features attributed to each node are constructed utilizing the worker’s personal profile information in conjunction with their task-related performance metrics.
Questions [28]: The dataset is derived from Yandex Q, a question-answering platform. The nodes represent individual users. An edge is established between two nodes to denote the scenario wherein one user has responded to another’s inquiry within the temporal bounds of one year, stretching from September 2021 to August 2022. This dataset is specifically curated to encapsulate the engagement of users who exhibit interest in the medical topic. The target for this dataset is to predict the likelihood of users’ continued activity on the platform by the conclusion of the observed period. The average FastText embeddings of the lexical content present in the users’ descriptions are used for node features.
Elliptic [10]: The dataset comprises a graph of 203,769 nodes representing Bitcoin transactions, connected through 234,355 edges representing payment flows, along with 166 distinctive node attributes. It correlates Bitcoin transactional data with corresponding real-world entities that are categorized as lawful, including exchanges, wallet services, mining operations, and legitimate services, as well as unlawful categories, including scams, malicious software, terrorist-related organizations, ransomware operations, and fraudulent investment activities known as Ponzi schemes.
DGraph [7]: The DGraph dataset is a large-scale graph with different edge types and temporal information supplied by Finvolution Group. It includes 3 million nodes, 4 million dynamic edges, and 1 million labeled nodes. The nodes represent user accounts within a financial organization offering personal loan services, while an edge between two nodes indicates that one account has designated the other as an emergency contact. Nodes classified as fraud correspond to users displaying delinquent financial conduct. For those accounts with borrowing records, outliers are identified as accounts with a history of overdue payments, while inliers are those without such a history. Additionally, the dataset includes 17 node features derived from user profile information.
IV-A2 Baselines
In our experiments, we evaluate GODM against different types of baseline methods. We have three types of baseline methods, including general graph neural networks, graph outlier detectors, and data augmentation for graph outlier detection. For general graph neural networks, we follow GADBench and select six GNNs:
GCN (Graph Convolutional Networks) [1]: GCN is the seminal work that applies convolution operation on graph data. It propagates the information of a node to its neighbors, thus enabling the network to develop a representation for each node that reflects its local neighborhood context.
SGC (Simplified Graph Convolution) [29]: This variant of GCN leverages Chebyshev polynomials to approximate the spectral graph convolution operator. This strategy allows the model to encompass both local and global graph structures, enhancing its scalability for handling larger graphs.
GIN (Graph Isomorphism Network) [30]: GIN is a form of GNN that effectively captures graph structures while maintaining graph isomorphism. It achieves this by generating consistent embeddings for structurally identical graphs, which is permutation invariant.
GraphSAGE (Graph Sample and Aggregate) [17]: GraphSAGE presents a general inductive learning approach where node embeddings are generated through the sampling and aggregation of features from a node’s immediate neighborhood.
GAT (Graph Attention Networks) [31]: GAT incorporates the attention mechanism within the GNN framework. It dynamically assigns varying importance to different nodes during the information aggregation process, focusing the model’s learning on the most relevant parts of the neighborhood.
GT (Graph Transformer) [32]: Drawing inspiration from the Transformer model in neural networks, GT adapts these principles for graph-structured data. It utilizes masks in the self-attention mechanism to capitalize on the inherent structure of graphs, thus boosting the model’s efficiency.
In addition to general GNNs, we also include eight GNNs specifically designed for graph outlier detection:
GAS (GCN-based Anti-Spam) [33]: GAS is an attention-based spam review detector, extending the capabilities of GCN to process heterogeneous and heterophilic graphs. It employs the KNN algorithm to align with the structure of each graph.
DCI (Deep Cluster Infomax) [34]: DCI is a self-supervised learning strategy that separates the learning of node representations from outlier detection. It addresses discrepancies between node behavioral patterns and their label semantics by clustering the entire graph, thus capturing intrinsic graph properties in focused feature spaces.
PCGNN (Pick and Choose Graph Neural Network) [35]: Tailored for imbalanced GNN learning in fraud detection scenarios, PCGNN uses a label-balanced sampler for node and edge selection during training. This results in a more balanced label distribution within the induced subgraph.
BernNet [36]: BernNet is a GNN variant offering a robust approach to designing and learning arbitrary graph spectral filters. It utilizes an order-K Bernstein polynomial approximation for estimating filters over the normalized Laplacian spectrum, catering to a variety of graph structures.
GATSep [37]: Designed to optimize learning on heterophily graphs, GAT-sep merges key design elements like ego- and neighbor-embedding separation, higher-order neighborhood processing, and combinations of intermediate representations.
AMNet (Adaptive Multi-frequency GNN) [38]: AMNet is structured to capture signals across both low and high frequencies by stacking multiple BernNets and adaptively integrating signals from different frequency bands.
BWGNN (Beta Wavelet Graph Neural Network) [39]: BWGNN addresses the ”right-shift” phenomenon on outliers. It uses the Beta kernel to address higher frequency anomalies through spatially/spectrally localized band-pass filters.
GHRN (Graph Heterophily Reduction Network) [40]: GHRN tackles the issue of heterophily in the spectral domain for graph outlier detection. This approach focuses on pruning inter-class edges to enhance the representation of high-frequency components in the graph spectrum.
For heterogeneous graphs, despite the rich literature, limited heterogeneous GNNs can cope with temporal information. As one of the representatives, we chose HGT for our baseline:
HGT (Heterogeneous Graph Transformer) [21]: HGT is engineered to address the challenges of modeling heterogeneous graphs. It introduces node- and edge-type dependent parameters that are crucial for defining heterogeneous attention mechanisms for each edge.
Because the data augmentation for graph outlier detection is a relatively new research topic, there are limited baselines. We include one baseline for homogeneous graphs and introduce a variant of GODM that can work on both homogeneous and heterogeneous datasets, GOVAE:
DAGAD (Data Augmentation for Graph Anomaly Detection) [41]: DAGAD incorporates three modules to augment the graph data, including an information fusion module for representation learning, a data augmentation module to enrich the training set with synthetic samples, and an imbalance-tailored learning module to distinguish between minority anomalous class and majority normal class.
GOVAE: In this variant, we drop Latent Diffusion Model and form a variational autoencoder with Variational Encoder and Graph Generator. The inference is a direct one-step estimation from prior Gaussian distribution.
Along with GODM, we compare these three data augmentation methods for graph outlier detection in Table II, where the flexibility represents whether the data augmentation is model agnostic. Heterogeneous and time indicate whether the method supports these two types of information. The diffusion column indicates whether the method adopts diffusion models.
TABLE II: Comparison of data augmentation for graph outlier detection.
Flexibility Heterogeneous Time Diffusion
DAGAD  ✗  ✗  ✗  ✗
GOVAE  ✗
GODM
IV-A3 Metrics
We follow the extensive literature in graph outlier detection [5, 6] to comprehensively evaluate the performance of graph outlier detectors with three metrics that are robust to class imbalance: Receiver Operating Characteristic-Area Under Curve (AUC), Average Precision (AP), and Recall@k (Rec), where the value of k is set to the number of actual outliers present in the dataset.
IV-A4 Implementation Details
We modified GADBench to benchmark outlier detection performance.
Environment. The key libraries and their versions used in our experiments are as follows: Python 3.9, CUDA 11.8, PyTorch 2.0.1 [42], PyTorch Geometric 2.4.0 [43], DGL 1.1.2 [44], and PyGOD 1.0.0 [4].
TABLE III: Performance comparison in AUC, AP, and Rec (%) on four datasets.
Dataset Weibo Tolokers Questions Elliptic
Metric AUC AP Rec AUC AP Rec AUC AP Rec AUC AP Rec
GCN 98.11 93.48 89.34 74.69 42.88 42.06 69.81 12.54 16.99 82.68 22.23 27.61
SGC 98.66 92.46 87.90 70.67 38.03 35.98 69.88 10.13 15.62 73.02 11.44 9.14
GIN 97.47 92.67 87.90 74.05 36.57 36.76 67.76 12.30 18.36 84.38 29.66 35.64
GraphSAGE 96.54 89.25 86.17 79.42 48.65 46.42 71.69 17.63 21.10 85.31 37.52 36.20
GAT 94.08 90.25 86.74 77.26 43.14 43.30 70.33 14.51 17.26 84.42 23.43 27.42
GT 97.06 91.44 87.03 79.24 46.22 46.57 70.83 16.14 20.27 87.14 29.91 38.97
[…]
DCI 93.90 87.78 83.86 75.98 39.85 40.19 67.95 14.58 19.18 81.93 27.63 33.15
[…]
GATSep 96.72 91.55 89.05 79.63 46.08 46.73 69.96 15.98 19.18 83.89 21.46 21.15
BernNet 93.85 88.00 85.30 76.20 42.20 42.21 70.80 16.04 17.53 82.01 20.52 23.55
AMNet 95.88 89.74 85.59 75.83 42.66 41.90 69.71 17.02 19.18 80.06 16.73 17.17
BWGNN 98.29 92.72 84.73 80.15 49.65 47.35 69.47 16.24 18.63 84.32 22.56 26.50
GHRN 97.21 92.67 88.18 79.80 49.50 48.29 68.24 16.24 18.63 85.36 24.01 30.29
DAGAD 98.54 83.36 90.78 77.69 33.94 44.39 71.21 6.88 20.55 85.62 26.18 40.54
[…]
Hardware. All of our experiments were performed on a Linux server with an AMD EPYC 7763 64-core CPU, 192GB RAM, and an NVIDIA RTX A40 GPU with 48GB memory.
Hyperparameters. GODM is implemented with default hyperparameters. For Variational Encoder, we use one layer GraphSAGE for $\text{GNN}_{\text{shared}}$, $\text{GNN}_{\bm{\mu}}$, and $\text{GNN}_{\bm{\sigma}}$, respectively. We set the hidden dimension to the largest power of 2, which is no greater than the feature dimension of the dataset divided by 2. The weights of reconstruction loss $\omega_{\mathbf{X}}$, $\omega_{\mathcal{E}}$, $\omega_{\mathbf{t}}$, and $\omega_{\mathbf{p}}$ are 1, 0.5, 1, and 0.3, respectively. The weight for KL-divergence $\beta$ is 0.001. For the diffusion model, we use the dimension twice as the hidden dimension in VAE. We use five layers of MLP as the denoising function following the detailed design in [22]. For training, we adopt the Adam optimizer with a learning rate of 0.001 without weight decay to train the VAE and the diffusion model for 100 epochs, respectively, and apply early stopping with a patience of 50. The negative sampling ratio is 1, and the approximate graph partition size is 2048. For inference, we use 50 steps of diffusion and generate the same amount of synthetic outliers as the number of real organic outliers in the training set. For all the hyperparameters in graph outlier detection, we apply the default setting in GADBench. We report the graph outlier detection performance of GODM with the optimal downstream graph outlier detector.
Open-Source Package. To enhance accessibility, we encapsulate our data augmentation into user-friendly API and make our code a plug-and-play package. This package is built upon PyTorch and PyTorch Geometric (PyG) frameworks. It accepts a PyG data object as input and returns the augmented graph as output. By inheriting the transforms.BaseTransform in PyG, we enable users to apply GODM just like any other transformations supported in PyG, e.g., ToUndirected. We are also able to customize hyperparameters during initialization. Code Demo 1 gives a brief example of API usage.
1from pygod.utils import load_data
2data = load_data(’weibo’) # load data
3
4from godm import GODM # import GODM
5godm = GODM(lr=0.004) # init. GODM
6aug_data = godm(data) # augment data
7
8detector(aug_data) # train on data
Code Demo 1: Using GODM on Weibo dataset [27].
For distribution, we release our package on the Python Package Index (PyPI). Users can easily install GODM by executing pip install godm in a command line.
IV-B Performamce on Graph Outlier Detection
TABLE IV: GODM’s performance improvements on graph outlier detection algorithms in AUC, AP, and Recall (%) on Tolokers.
AUC AP Rec
GCN 75.45 (+0.76) 44.17 (+1.29) 44.24 (+2.18)
SGC 72.73 (+2.06) 39.75 (+1.72) 38.01 (+2.02)
GIN 74.83 (+0.78) 38.54 (+1.96) 38.32 (+1.56)
GraphSAGE 81.65 (+2.23) 52.53 (+3.87) 50.93 (+4.52)
GAT 82.18 (+4.91) 51.13 (+7.99) 50.78 (+7.48)
GT 82.73 (+3.49) 51.73 (+5.51) 50.93 (+4.36)
GAS 76.96 (+0.05) 45.48 (-1.87) 43.15 (-1.87)
DCI 73.75 (-2.23) 37.52 (-2.33) 37.54 (-2.65)
PCGNN 73.65 (+1.47) 38.42 (+0.90) 38.01 (+1.25)
GATSep 83.46 (+3.83) 52.95 (+6.87) 52.80 (+6.07)
BernNet 76.53 (+0.33) 44.04 (+1.83) 42.21 (+0.00)
AMNet 76.67 (+0.84) 44.31 (+1.65) 43.30 (+1.40)
BWGNN 82.10 (+1.95) 51.94 (+2.29) 51.25 (+3.89)
GHRN 82.05 (+2.25) 51.84 (+2.34) 51.56 (+3.27)
We start by examining how effective GODM is in the task of graph outlier detection on homogeneous graphs. We report graph outlier detection performance on four homogeneous datasets in terms of AUC, AP, and Rec of different algorithms in Table III. In the table, the highest score in each metric is marked in bold, while the second highest score is underlined. For DAGAD, we report the optimal performance of its two implementations with GCN and GAT. Although DAGAD shows superior performance compared to the standalone GCN and GAT, it fails to surpass more sophisticated detectors. On the other hand, GODM and GOVAE together demonstrate significant enhancements in graph outlier detection performance across all datasets and metrics. Notably, on the Weibo dataset, where the performance metric is already relatively high (AUC of 98.66 for SGC), GODM further elevates the performance, achieving an AUC of 99.57. A particularly noteworthy improvement is observed on Elliptic, where GODM enhances the Rec by over 20% (an increase of 9.14) compared to GAS, which is the second-best performing algorithm. Although GOVAE exhibits competitive performance, its limited generation ability results in generally weaker performance compared to GODM, especially on Elliptic.
Further investigations were carried out to understand the extent of improvement GODM could bring to various graph outlier detection algorithms, specifically on Tolokers. We integrate GODM on the top of all graph outlier detection algorithms and report the performance changes in terms of AUC, AP, and Rec in Table IV. The absolute performance is reported out of brackets, and the relative change is reported in the brackets. We can observe that GODM enhances graph outlier detection performance for most of the algorithms. GODM increases the AP of GAT by 7.99. For DCI and GAS, the only two algorithms that minorly decrease the performance, this could be attributed to suboptimal hyperparameter settings for these specific algorithms or the dataset. We believe the performance can be improved by careful hyperparameter tuning.
TABLE V: Performance comparison in AUC, AP, and Rec (%) on DGraph.
AUC AP Rec
HGT 72.89 3.13 5.20
w/o Feature 75.10 3.64 5.59
w/o Edge 74.36 3.52 5.76
w/o Time 74.23 3.29 4.73
w/o Type 75.09 3.62 5.98
GOVAE 75.36 3.58 5.54
GODM 75.80 3.66 6.19
Moreover, in Table V, we benchmark the graph outlier detector performance on DGraph, a large-scale dataset with different types of edges and temporal information. We include HGT along with variants of GODM. For the w/o Feature, we randomly sample the node feature from the Gaussian distribution to replace the generated node feature in the synthetic graph. For the w/o Edge, all generated edges are substituted with random edges. For w/o Time, we give uniformly sampled timestamps to replace the generated timestamps. For the w/o Type, we use a random type sampled from a uniform distribution for all generated edges. According to Table V, GODM demonstrates enhanced performance across all metrics when compared to HGT. Furthermore, masking any generated information leads to a notable decline in performance, which underscores the significance of each aspect of the generated data. While GOVAE achieves commendable results, its performance still falls short of GODM’s, highlighting the superior efficacy of Latent Diffusion Model.
IV-C Generation Quality
To assess the quality of generated data, we conduct a case study on DGraph to compare real data and synthetic data from both GOVAE and GODM. The results are shown in Fig. 3. Specifically, Fig. 3 (a) and (b) illustrate the distribution density of a single dimension in node features. From the figures, we can see that the distributions of GODM generated node features are close to the complex distributions of real data, while GOVAE can only generate Gaussian distributions, showcasing the generation ability of Latent Diffusion Model. In addition, to evaluate generated edges, Fig. 3 (c) shows the frequency across eleven different edge types. The edge type generated by GODM has a more similar distribution to real data compared to GOVAE, particularly on the 10th edge type. It further underscores the superiority of GODM over GOVAE.
Figure 3: Visualization of single node feature density and edge type frequency of real data and synthetic data from GOVAE and GODM on DGraph.
IV-D Efficiency Study
We further evaluate the efficiency of GODM and its variants in terms of time and memory on graphs of various sizes, including Weibo, Elliptic, and DGraph. In order to demonstrate the efficiency of GODM, we remove graph clustering (w/o GC), negative sampling (w/o NS), and EDM (w/o EDM) for evaluation. To simulate the real-world application, we measure the running time of unit epoch training for VAE and Latent Diffusion Model plus the inference process. For memory, we present the maximum active GPU memory usage within the whole data augmentation, as the GPU memory constraint is usually the bottleneck for machine learning systems. The results are shown in Fig. 4.
Figure 4: Running time and GPU memory of GODM and its variants.
From Fig. 4, we can observe that employing a graph clustering algorithm significantly diminishes GPU memory utilization. Specifically, on DGraph, w/o GC results in a memory usage that is over 30$\times$ higher than GODM. On the other hand, graph clustering turns full-batch training into mini-batch training, leading to an anticipated increase in running time. Additionally, negative sampling reduces both the running time and the memory usage, as the number of edges in the training set reduces from $O(n^{2})$ to $O(|\mathcal{E}|)$. While EDM maintains consistent memory consumption, it contributes to time savings by minimizing the number of sampling steps.
V Related Works
This section briefly summarizes the previous related works in three key areas, including graph outlier detection, class imbalance, and graph generative models.
V-A Graph Outlier Detection
Graph outlier detection is a vital task in data mining and machine learning, aiming to identify anomalous structures (e.g., nodes, edges, subgraphs) in graphs. Due to its structural simplicity and broader applications, most research focuses on node-level graph outlier detection and can easily convert to edge level and graph level task [4]. Many studies focus on unsupervised settings, detecting outliers in the graph only based on the graph data itself without any ground truth labels [45, 46, 47]. However, such unsupervised approaches may not align well with scenarios necessitating the detection of specific outlier types containing domain-specific knowledge. In this case, (semi-)supervised graph outlier detectors, which can learn ground truth labels, are better options. [33] apply attention mechanism to detect spam review. [11, 48] alleviate the camouflage issue in fraud detection by message passing in selected neighborhoods. [35] further improves the neighbor selection in fraud detection. [49] builds the outlier detection graph via the KNN algorithm. [34] decouples the representation learning and anomaly detection, and [50] focuses on heterophily in graph outlier detection. Some other works detect outliers from a spectral perspective. [36] learns arbitrary graph spectral filters via Bernstein approximation, and [38] is capable of discerning both low-frequency and high-frequency signals, thereby adaptively integrating signals across a spectrum of frequencies. In [39], the Beta kernel is employed to detect anomalies at higher frequencies using flexible and localized band-pass filters. [40] addresses heterophily in graph outlier detection from a graph spectrum perspective.
V-B Class Imbalance
Despite the fruitful literature on graph outlier detection, most of methods underrate a pivotal challenge: the class imbalance in ground truth labels. These methods typically mitigate this imbalance by primitive techniques. GADBench uses reweighting [6], assigning a higher weight to outliers in the loss function. Some works adopt upsampling to replicate outliers, which is mathematically equivalent to reweighting. [11] applies downsampling, reducing the number of normal instances and losing a large amount of valuable supervision information from normal instances. However, these tricks may not sufficiently address the underlying complexities associated with imbalanced datasets. Consequently, there is a compelling need for more sophisticated techniques that can effectively address this class imbalance to ensure the quality of graph outlier detection. [35] balances the neighborhood distribution by reducing edges between different classes and adding edges between the same classes but does not change the overall number of positive and negative nodes. [41] adopts random perturbation for data augmentation to generate more samples. Other general graph machine learning studies apply interpolation-based data augmentation methods in latent space [51]. However, these methods are not specifically designed for graph outlier detection, and interpolation is too naive to generate helpful instances. This paper seeks to build upon these foundational works, proposing a novel graph data augmentation algorithm with latent diffusion models that not only mitigate class imbalance effectively but are also scalable and efficient on large graphs, ensuring broader applicability in real-world graph outlier detection.
V-C Graph Generative Models
Recent advances in graph generative models have catalyzed a significant body of research, focusing primarily on the synthesis of realistic and structurally coherent graphs. [52] proposes VGAE, leveraging variational autoencoder frameworks to learn latent representations of graph structures for generation purposes. [53] introduces GraphRNN, generating nodes and edges autoregressively with two recurrent neural networks. [54] applies a generative adversarial networks framework to generate graphs via random walks. [55] introduces the normalizing flow method for molecular graph generation. With the recent proliferation of works in image generation with diffusion models, some efforts have extended the generation power of diffusion models to graph-structured data. [13] proposes a score-based generative modeling of graphs via the system of stochastic differential equations, while [14] uses a discrete diffusion model to generate graphs. However, the potential of graph generative models to address the class imbalance in graph outlier detection remains largely underexplored. We propose to generate some synthetic outliers to mitigate this class imbalance in graph outlier detection.
VI Conclusion
In this paper, we introduce GODM, a novel method for data augmentation in supervised graph outlier detection. GODM is designed to address the significant challenge of class imbalance, a pervasive issue in graph outlier detection, where the scarcity of outlier instances compared to normal instances often leads to suboptimal model performance. Our method leverages the power of generative models, particularly latent diffusion models, to synthesize fidelity graph data that is statistically similar to real outliers. GODM consists of three key components: Variational Encoder, Latent Diffusion Model, and Graph Generator. Variational Encoder maps heterogeneous information in the input graph into a unified latent space. Latent Diffusion Model learns the latent space distribution of real organic data through iterative denoising. Graph Generator then synthesizes graph data from this latent space embedding, ensuring statistical similarity to real data. Our extensive experiments across multiple datasets demonstrate the effectiveness and efficiency of GODM. Note that GODM is model agnostic, which means it can be flexibly integrated with different downstream graph outlier detectors. We encapsulate GODM into a plug-and-play package, making it accessible for broader use in the researcher and practitioner communities.
In future research, several directions can be explored to enhance GODM further and expand its applicability in the domain of graph outlier detection:
  • Diffusion in graph sapce. Considering the heterogeneity and scalability of the graphs for outlier detection applications, GODM currently employs a diffusion model in latent space. [56] starts to generate large graphs directly in graph space. Future research could explore the application of diffusion models on heterogeneous information directly in graph space, potentially enhancing the model’s ability to capture complex graph structures.
  • Collaboration with outlier detection. Currently, GODM is model agnostic, meaning the training of GODM and downstream outlier detection tasks are separate. Integrating the supervision signals from both the generation task and the downstream outlier detection in a unified training framework could be mutually beneficial. Self-training offers a promising paradigm for this integration [57].
  • Enhancing generative models. Although the current diffusion model in GODM has shown impressive capabilities, there is potential for further improvement. This could be achieved by employing more expressive denoising functions, such as transformers [58], or by exploring alternative generative models like energy-based models [59] and or normalizing flows [60], which might offer different advantages over diffusion models.
Acknowledgment
This work is supported by NSF under grant III-2106758. We sincerely appreciate reviewers in advance for their significant effort and invaluable feedback during the review process.